# milliFlow: Scene Flow Estimation on mmWave Radar Point Cloud for Human Motion Sensing

### 12 Jan 2024 v4

@@ Abstract.
Approaching the era of ubiquitous computing, human motion sensing plays a crucial role in smart systems for decision making, user interaction, and personalized services. Extensive research has been conducted on human tracking, pose estimation, gesture recognition, and activity recognition, which are predominantly based on cameras in traditional methods. However, the intrusive nature of cameras limits their use in smart home applications. To address this, mmWave radars have gained popularity due to their privacy-friendly features. In this work, we propose milliFlow, a novel deep learning method for scene flow estimation as a complementary motion information for mmWave point cloud, serving as an intermediate level of features and directly benefiting downstream human motion sensing tasks. Experimental results demonstrate the superior performance of our method with an average 3D endpoint error of 4.6cm, significantly surpassing the competing approaches. Furthermore, by incorporating scene flow information, we achieve remarkable improvements in human activity recognition, human parsing, and human body part tracking. To foster further research in this area, we will provide our codebase and dataset for open access upon acceptance.
Scene Flow Estimation, Radar Point Cloud, mmWave Human Motion Sensing.
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXjournal: IMWUTjournalvolume: 0journalnumber: 0publicationmonth: 11ccs: Human-centered computing → Ubiquitous and mobile computingccs: Computer systems organization → Embedded and cyber-physical systems
1. Introduction
@@
Perceiving and understanding human behaviours play a pivotal role in human-centred applications such as disaster response (Tariq et al., 2018; Mishra et al., 2020), surveillance (Paul et al., 2013; Baltieri et al., 2011) and health monitoring (Mukhopadhyay, 2014; Del Rosario et al., 2015; Fagert et al., 2019). Conventional methods rely on cameras (Güler et al., 2018; Shu et al., 2020) or wearables (Lemmens et al., 2015; Wang et al., 2018; Bannis et al., 2023) which are prone to visual deterioration (such as low lighting conditions, smoke, and fog) and raise privacy concerns, potentially compromising user experience with intrusive measures. To address these concerns, researchers, on the other side, also propose to use wireless radio frequency (RF) signals bounced off the human body for human sensing (Tian et al., 2018; Kianoush et al., 2016; Zhang et al., 2016, 2020, 2019) which is robust against poor lighting, privacy-preserving and non-intrusive to users. Among many RF techniques, single-chip millimetre wave (mmWave) radar emerges as a low-cost sensor that can provide more trustworthy point clouds of a scene under environment dynamics due to its transceiver design and the usage of Continuous Wave (CW). For these reasons, there has been a significant increase in the production of single-chip radars (iwr, 2023) and wide deployment in real-world scenarios, ranging from smart buildings (who, 2023; Tseng et al., 2014) to vehicle cabins (Schwarz et al., 2021; Chen et al., 2021), and to first-responder toolkits (res, 2023).
@@
Figure 1. We propose milliFlow, a scene flow estimation module to provide an additional layer of point-wise motion information on top of the original mmWave radar point cloud in the conventional radar human motion sensing pipeline.
@@
As motion is naturally continuous and human bodies are non-rigid, point-wise velocity per radar frame is intuitively a strong cue for improving the motion estimation robustness and has been widely used in prior arts (Xue et al., 2021; Kong et al., 2022; Wang et al., 2023; Singh et al., 2019). However, fine-grained velocity is difficult to obtain when it comes to mmWave radars. First, while some radars provide the Doppler velocity of a point, the velocity resolution is rather low (Iovescu and Rao, 2017) and thus cannot accurately capture the subtle human body movement, which is usually slower than 0.5m/s in domestic life. Moreover, mmWave radar only senses radial Doppler information but fails to capture the tangential one. For some radars designed for human sensing, e.g., the Vayyar radar (vay, 2023) used in this work, the Doppler information is even absent in their sensor outputs. Second, extracting motion information from consecutive mmWave radar point clouds is also highly error-prone due to the limitations inherent in low-cost single-chip mmWave radar. These limitations include extremely sparse point clouds due to the target detector, e.g., constant false alarm rate (CFAR) algorithm (Scharf and Demeure, 1991) used on-chip and the presence of ghost points caused by multi-path effect (Gennarelli and Soldovieri, 2014). More recently, it has been found that in a single radar frame, only a subset of body parts reflecting the signal towards the radar can be observed, while other parts deflecting the signal away from the radar are missing from the capture (Cao et al., 2022). As a result, some human body parts detected in one frame may disappear in the next frame. For the above reasons, conventional radar point tracking methods (Rohal and Ochodnicky, 2017; Yan et al., 2022; Godrich et al., 2010) struggle to track human motion across frames or give fine-grained velocity in between.
@@
In this work, we propose to estimate and use scene flow as intermediate features to better support radar-based human sensing. Scene flow refers to a set of displacement vectors between two consecutive point-cloud frames describing the motion field of a 3D scene. We hypothesise that scene flow, if estimated accurately, is able to drastically facilitate cross-frame movement analysis by directly exposing per-point motion features, thereby addressing the aforementioned challenges in previous research. Scene flow has long been proven very effective by the computer vision community in image-based human motion sensing applications (Wang et al., 2017; Liu et al., 2012; Jaimez et al., 2015). However, estimating scene flow on mmWave radar point cloud is non-trivial because of the inherent sparsity and noise of radar point cloud. Directly applying conventional scene flow estimation methods designed for LiDAR or RGB-D cameras (Wei et al., 2021; Wu et al., 2020; Cheng and Ko, 2022; Kittenplon et al., 2021) on radar point clouds has been found inadequate and generalises poorly. On the other side, recent mmWave radar scene flow estimation works (Ding et al., 2022, 2023) generally focus on autonomous driving scenarios and rely on deep learning-based pipelines. However, these existing methods cannot be readily applied to our human sensing scenarios because the rigid-body assumption used for autonomous driving scenarios cannot stand in our cases where human subjects have non-rigid motion.
@@
To cope with the above challenges, we propose milliFlow, a novel mmWave radar-based scene flow estimation approach for human motion sensing scenarios. Our contributions include:
@@
  • To the best of our knowledge, milliFlow is the first-of-its-kind work that aims to estimate mmWave radar-based scene flow for human motion sensing applications.
@@
  • We address the technical challenges, e.g., sparsity, and lack of temporal cues, for scene flow estimation learning in our cases with a bespoken end-to-end learning network.
@@
  • We collect a large-scale human sensing dataset with 12 participants performing 8 different activities in three sites for evaluation, and perform a comprehensive evaluation of scene flow estimation accuracy as well as the performance on three downstream tasks.
@@
  • We will open source our code and dataset to benefit the mmWave radar human sensing research communities.
2. SFCW mmWave radar preliminary
In this work, we employ the Vayyar radar (vay, 2023), which is a stepped-frequency continuous-wave (SFCW) radar (Weiss, 2009) bespoken designed for fine-grained human sensing. This section aims to provide a concise overview of the point cloud generation principles of SFCW radar to aid in comprehending the hardware components of the proposed system. In what follows we introduce the working principle of SFCW radars.
(1) $\Delta\phi=\phi_{2}-\phi_{1}=\frac{4\pi R}{c}(f_{2}-f_{1})=\frac{4\pi R}{c}\Delta f$
and we get
(2) $R=\frac{c}{4\pi\Delta f}\Delta\phi$
Also, we have $\Delta\phi\in[0,2\pi]$, so the maximum unambiguous range is inversely proportional to $\Delta f$, which is the frequency step in SFCW radar. During each frame, the transmitter sequentially emits a series of waves of increasing frequency at equal intervals. The transmitted signal is reflected back at different surfaces in the scene and received by the receiving antenna. The ADC samples the change in amplitude and phase between the transmitted signal and the received signal and stores the values in IQ format. We can apply Fourier Transform to get the phase variation, with which we can further derive the range of the objects. This is also known as ”Range-FFT”. The range resolution is inversely proportional to the bandwidth of the radar, written as $R_{res}=\frac{c}{2B}$, where $B=f_{max}-f_{min}$.
Angle-of-arrival Estimation. Angle-of-Arrival (AoA) estimation can be performed using a linear receiver array with equally-spaced elements, typically separated by a distance of $1/2\lambda$, where the number of elements in the array, denoted as $N$, is greater than or equal to two. When a reflected signal is received at each element of the array, a phase difference arises due to the slight differences in signal travel distance, as determined by the angle of arrival, $\theta$. This phase difference can be quantified using the formula $\Delta\phi=kdsin(\theta)$, where $k$ represents the wave number and $d$ is the distance between consecutive elements. The angle of arrival can then be estimated by analyzing the phase difference using a Fourier Transform technique, commonly referred to as ”Angle-FFT”. Vayyar uses Multiple-Input Multiple-Output (MIMO) array, and for each frame, we are able to get a 2D virtual antenna matrix, from which we can estimate azimuth and elevation AoA simultaneously.
Point Cloud Generation. For each frame, the raw data is stored as a complex matrix of size $N_{a}*N_{e}*M$, where $N_{a}$ and $N_{e}$ are the numbers of Tx-Rx antenna pairs (i.e., virtual antennas) in azimuth and elevation directions, respectively, and M is the number of frequency steps. First, the clutter removal is applied to focus on dynamic objects. We then apply Range-FFT for Range estimation for each virtual antenna and further perform 2D Angle-FFT to get the AoA information in both directions. Following the above processing, we are able to get a 3D heatmap for each frame. Strong peaks are then detected with methods like CFAR (Scharf and Demeure, 1991), and converted to 3D points, which is the input to our scene flow estimation.
3. Technical Challenges
After understanding the working principles behind the SFCW mmWave radar, we now elaborate on the technical challenges of using it for scene flow estimation under human motion sensing scenarios.
Sparsity and Noise. Due to the bandwidth and hardware limitations, the heatmap used to detect targets (c.f. Sec. 2) has low resolution in both range and angular dimensions. Moreover, only outstanding peaks are selected from the heatmap as valid targets. As a result, the radar point cloud returned by related algorithms (e.g. CFAR (Scharf and Demeure, 1991)) is inevitably sparse with only an average of $\sim$100 points correlated to the human body. Some body parts even do not have their correlated radar points. More seriously, due to the multi-path effect of mmWave, there are a non-negligible number of ghost points in the radar point cloud that further poisons the original sparse radar data. Such sparse and noisy natures yield mmWave radar scene flow estimation non-trivial as the local geometric cues tend to be insufficient and low-quality. Concretely, it is hard to extract strong local features for each radar point with their scarce or absent neighbour points, which is further exacerbated by the non-negligible noisy points in radar point clouds, making the local features less robust.
Lack of Temporal Cues. Representing the radial velocity of each point, the radar Doppler velocity measurement has great potential to benefit our radar scene flow estimation. Unluckily, the resolution of Doppler velocity measurement is limited by the hardware and is usually inadequate to identify the real motion status of objects, especially in human sensing scenarios, where the scale of movement is small. Furthermore, some radar products, for example, the Vayyar radar (vay, 2023) used in this work, do not support the measurement of the Doppler velocity with only one excerpt of CW emitted per frame. The low resolution or absence of the Doppler velocity makes accurate scene flow estimation difficult. Particularly for those points that also do not have associated points in the next frame, they cannot get any useful guidance from both two worlds for scene flow estimation. Unfortunately, certain body parts (e.g. head, arm and leg) are often missing in radar point clouds, which means they do not have correlated radar points in the vicinity. Such inconsistent local capture together with the issue of Doppler velocity jointly pose another challenge to our method in the aspect of temporal information.
Scene Flow Annotation. In this work, as we rely on deep networks to learn accurate scene flow estimation, it is essential for us to have point-level scene flow labels that can be leveraged to supervise the network output. However, manually annotating point-wise scene flow is extremely expensive and suffers from the lack of real correspondence between real-world point clouds. Therefore, instead of labelling point by point from the ground, some recent works (Jund et al., 2021; Huang, Shengyu and Gojcic, Zan and Huang, Jiahui and Wieser, Andreas and Schindler, Konrad, 2022; Ding et al., 2023) adopt a top-down scheme that only annotates the bounding box of ambient objects and utilizes them to generate point-level pseudo scene flow ground truth. This object-based labelling scheme works well for autonomous driving scenarios as almost all road users (e.g. car, truck, motorcycle) can be seen as rigid bodies. However, despite the cost used for obtaining object annotations, the rigid-body assumption cannot hold in our human sensing applications where the motion of each human subject is non-rigid.
4. Methodology
We now introduce our proposed method to address the above technical challenges.
4.1. Overview
We formally formulate the scene flow estimation problem in Sec. 4.2. Sec. 4.3 details the design of the neural network that takes consecutive radar point clouds as the input and outputs the scene flow in between. Particularly, to address the sparsity and noise challenges, we propose to aggregate the global features to complement the vulnerable local features (c.f. Sec. 4.3.2). In the form of representation vectors, such global features can aggregate the features of all points and provide holistic information about a point cloud. Then, to compensate for the absence of Doppler velocity and inconsistent local capture, we propagate the temporal information to the current frame using a GRU network (Cho et al., 2014) to update the global vectors of the flow embedding (c.f. Sec. 4.3.4). This strategy is intuitive as radar data is captured in sequence. Fig. 2 illustrates the architecture of our scene flow network. Last but not least, to efficiently label point-based scene flow for training, we propose a cross-modal automatic scene flow labelling scheme (c.f. Sec. 4.4), specific to point clouds captured in human sensing scenarios. Our motivation is based on the observation that the human body can be roughly segmented into multiple skeletons each of which can be seen as rigid bodies, such as the upper arm, neck and thigh bone. We then attribute the human dynamics to the rigid motion of these skeletons and further assume that the scene flow of points in the vicinity is induced by them. Given this assumption, we can first calculate the rigid transformation of skeletons and use them to annotate scene flow vectors for points attributed to them. The labelling pipeline can be seen in Fig. 3.
4.2. Problem Formulation
@@
Here, we consider the problem of scene flow estimation for dynamic 3D point clouds collected by an SFCW mmWave radar sensor (c.f. Section 2) in human sensing scenarios. As a general problem setting, the input to point cloud-based scene flow estimation is two consecutive 3D point clouds $\mathcal{P}=\{p_{i}\}^{N}_{i=1}$ and $\mathcal{Q}=\{q_{i}\}^{M}_{i=1}$ captured by the same device and the output is a set of 3D vectors =$\mathcal{F}=\{f_{i}\}^{N}_{i=1}$ that align each point ${p}_{i}$ in $\mathcal{P}$ to its associated position ${p}^{a}_{i}={p}_{i}+f_{i}$ in the frame of $\mathcal{Q}$. Different from the task that finds real correspondences between two frames, scene flow estimation only derives per-point 3D displacement for $\mathcal{P}$ and the associate position ${p}^{a}_{i}$ does not necessarily overlap with any points in $\mathcal{Q}$. Besides 3D coordinates information, each point may also have additional properties given mmWave radar point clouds as input, such as Doppler velocity or intensity value. Without loss of generality, here we concatenate the per-point incident properties and 3D point coordinates into the 2D matrix and use $\mathcal{X}=\{x_{i}\}^{N}_{i=1}$, $\mathcal{Y}=\{y_{i}\}^{M}_{i=1}$ to denote the data from the source and target frame, respectively.

### 13 Mar 2024 v5

@@ Abstract
Human motion sensing plays a crucial role in smart systems for decision-making, user interaction, and personalized services. Extensive research that has been conducted is predominantly based on cameras, whose intrusive nature limits their use in smart home applications. To address this, mmWave radars have gained popularity due to their privacy-friendly features. In this work, we propose milliFlow, a novel deep learning approach to estimate scene flow as complementary motion information for mmWave point cloud, serving as an intermediate level of features and directly benefiting downstream human motion sensing tasks. Experimental results demonstrate the superior performance of our method when compared with the competing approaches. Furthermore, by incorporating scene flow information, we achieve remarkable improvements in human activity recognition and human parsing and support human body part tracking. To foster further research in this area, we will provide our codebase and dataset for open access.
1 Introduction
@@
Perceiving and understanding human behaviours play a pivotal role in human-centred applications such as disaster response [68, 54], surveillance [58, 8] and health monitoring [56, 20, 25]. Conventional methods rely on cameras [31, 66] or wearables [45, 74, 9] which are prone to visual deterioration (such as low lighting conditions, smoke, and fog) and raise privacy concerns, potentially compromising user experience with measurements that are psychologically intrusive. To address these concerns, researchers, on the other side, also propose to use wireless radio frequency (RF) signals bounced off the human body for human sensing [71, 42, 87, 89, 88] which is robust against poor lighting, privacy-preserving and non-intrusive to users. Among many RF techniques, single-chip millimetre wave (mmWave) radar emerges as a low-cost sensor that can provide more trustworthy point clouds of a scene under environment dynamics due to its MIMO transceiver design. For these reasons, there has been a significant increase in the production of single-chip radars [2] and wide deployment in real-world scenarios, ranging from smart buildings [5, 72] to vehicle cabins [65, 16], and to first-responder toolkits [3].
@@
Figure 1: We propose milliFlow, a scene flow estimation module to provide an additional layer of point-wise motion information on top of the original mmWave radar point cloud in the conventional mmWave-based human motion sensing pipeline.
@@
As motion is naturally continuous and human bodies are non-rigid, point-wise velocity per radar frame is intuitively a strong cue for improving the motion estimation robustness and has been widely used in prior arts [85, 44, 77, 67]. However, fine-grained velocity is difficult to obtain when it comes to mmWave radars. First, while some radars provide the Doppler velocity of a point, the velocity resolution is rather low [36] and thus cannot accurately capture the subtle human body movement, which is usually slower than 0.5m/s in domestic life. Moreover, mmWave radar only senses radial Doppler information but fails to capture the tangential one. For some radars designed for human sensing, e.g., the Vayyar radar [4] used in this work, the Doppler information is even absent in their sensor outputs. Second, extracting motion information from consecutive mmWave radar point clouds is also highly error-prone due to the limitations inherent in low-cost single-chip mmWave radar. These limitations include extremely sparse point clouds due to the target detector, e.g., constant false alarm rate (CFAR) algorithm [64] used on-chip and the presence of ghost points caused by multi-path effect [26]. More recently, it has been found that in a single radar frame, only a subset of body parts reflecting the signal towards the radar can be observed, while other parts deflecting the signal away from the radar are missing from the capture [13]. As a result, some human body parts detected in one frame may disappear in the next frame. For the above reasons, conventional radar point tracking methods [63, 86, 27] struggle to track human motion across frames or give fine-grained velocity in between.
@@
In this work, we propose to estimate and use scene flow as intermediate features to better support radar-based human sensing. Scene flow refers to a set of displacement vectors between two consecutive point-cloud frames describing the motion field of a 3D scene. We hypothesise that scene flow, if estimated accurately, is able to drastically facilitate cross-frame movement analysis by directly exposing per-point motion features, thereby addressing the aforementioned challenges in previous research. Scene flow has long been proven very effective by the computer vision community in image-based human motion sensing applications [76, 51, 37]. However, estimating scene flow on mmWave radar point cloud is non-trivial because of the inherent sparsity and noise of radar point cloud. Directly applying conventional scene flow estimation methods designed for LiDAR or RGB-D cameras [81, 83, 17, 43] on radar point clouds has been found inadequate and generalises poorly. On the other side, recent mmWave radar scene flow estimation works [23, 22] generally focus on autonomous driving scenarios and rely on deep learning-based pipelines. However, these existing methods cannot be readily applied to our human sensing scenarios because the rigid-body assumption used for autonomous driving scenarios cannot stand in our cases where human subjects have non-rigid motion.
@@
To cope with the above problems, we propose milliFlow, as exhibited in Fig. 1, a novel mmWave radar-based scene flow estimation approach for human motion sensing scenarios. Our contributions include:
  • $\bullet$
@@
    To the best of our knowledge, milliFlow is the first-of-its-kind work that aims to estimate mmWave radar-based scene flow for human motion sensing.
  • $\bullet$
@@
    We address the challenges, e.g., sparsity, and lack of temporal cues, for scene flow learning in our cases with a bespoken end-to-end learning network.
  • $\bullet$
@@
    We collect a large-scale human motion sensing dataset for evaluation, and perform a comprehensive evaluation of scene flow estimation accuracy as well as the performance on three downstream tasks.
  • $\bullet$
@@
    We will open-source our code and data to benefit the mmWave radar human sensing research communities.
2 Related Work
mmWave Radar-based Human Sensing. The feasibility and versatility of mmWave radar have been extensively demonstrated in various human sensing applications, including vital sign monitoring [7, 78, 53], signature verification [32, 47], fall detection [38, 73], human tracking and identification [90, 91, 19, 15, 30], gesture recognition [50, 33, 57, 49], activity recognition [67, 79, 6] and pose estimation, reconstruction [85, 84, 44]. Compared with these applications, our work is unique in that we aim to estimate point-level scene flow vectors instead of providing a holistic output for the whole point cloud. In this way, we can either explicitly augment each radar target with scene flow vectors or implicitly learn robust latent features, which can further benefit many downstream tasks (cf. Sec. 4.4).
Scene Flow Estimation on Point Clouds. Recent scene flow works are mostly towards autonomous driving applications and attempt to estimate scene flow on LiDAR point clouds captured from autonomous vehicles. To this goal, different approaches are proposed, including classical methods [21, 59, 48] and deep learning-based ones [52, 10, 83, 81]. Recently, with the advances in point cloud feature learning, deep learning-based methods become more prevalent. According to their learning paradigm, prior works in this thread can be divided into fully-supervised [80, 75, 60, 81, 11], self-supervised [43, 46, 10, 55, 83] and weakly-supervised [28, 24] learning approaches. Apart from the aforementioned works, our work aims to estimate scene flow for human sensing using mmWave radar. Moreover, our method does not demand any manual annotation efforts, instead, we automatically generate noisy pseudo scene flow labels from corresponding RGB-D images captured by the co-located camera.
Scene Flow Estimation with mmWave Radar. As far as we know, there are limited works that estimate scene flow using mmWave radar. A pioneering work [23] introduced a self-supervised learning method tailored for automotive radar, utilizing unique loss functions and a two-stage network. To improve scene flow performance and enable more downstream applications, a later work [22] exploits cross-modal supervision from co-located sensors (e.g. IMU, LiDAR) on modern autonomous vehicles for radar scene flow learning. However, these methods cannot be transferred to human sensing scenarios due to two reasons. First, the radar used for human-centric applications is different from the automotive radar in many aspects. For example, automotive radar has a lower range but higher angular resolution, making scene flow models bespoken for them hard to generalize to human sensing. More importantly, the human object is a non-rigid body, while in autonomous driving, scene dynamics are usually attributed to rigid body motion (e.g. cars and motorcycles) [28, 21, 24]. Such discrepancy in objective indicates different label and supervision generation schemes.
3 Methodology
3.1 Overview
We formally formulate our scene flow estimation problem in Sec. 3.2. Then we elaborate on the technical challenges of using mmWave radar for scene flow estimation under human motion sensing scenarios in Sec. 3.3. Sec. 3.4 details the design of our neural network, which comprises five sequential modules, tackling the sparsity and noise challenges and compensating for the lack of temporal cues. In Sec. 3.5, we propose a cross-modal automatic scene flow labelling scheme, to efficiently label scene flow for point clouds captured in human sensing scenarios. Sec. 3.6 further introduces our loss function used for training scene flow network.
3.2 Problem Formulation
@@
Here, we consider the problem of scene flow estimation for dynamic 3D point clouds collected by an mmWave radar sensor used in human sensing scenarios. As a general problem setting, the input to point cloud-based scene flow estimation is two consecutive 3D point clouds $\mathcal{P}=\{p_{i}\}^{N}_{i=1}$ and $\mathcal{Q}=\{q_{i}\}^{M}_{i=1}$ captured by the same device and the output is a set of 3D vectors =$\mathcal{F}=\{f_{i}\}^{N}_{i=1}$ that align each point ${p}_{i}$ in $\mathcal{P}$ to its associated position ${p}^{a}_{i}={p}_{i}+f_{i}$ in the frame of $\mathcal{Q}$. Different from the task that finds real correspondences between two frames, scene flow estimation only derives per-point 3D displacement for $\mathcal{P}$ and the associate position ${p}^{a}_{i}$ does not essentially overlap with any points in $\mathcal{Q}$. Besides 3D coordinates information, each point may also have additional properties given mmWave radar point clouds as input, such as Doppler velocity or intensity value. Without loss of generality, here we concatenate the per-point incident properties and 3D point coordinates into the 2D matrix and use $\mathcal{X}=\{x_{i}\}^{N}_{i=1}$, $\mathcal{Y}=\{y_{i}\}^{M}_{i=1}$ to denote the data from the source and target frame, respectively.
3.3 Technical Challenges
Sparsity and Noise. Due to bandwidth and hardware limitations, the radar raw data has low resolution in both range and angular dimensions from which only outstanding peaks are selected as valid targets. As a result, the point cloud generated by mmWave radar is quite sparse with an average of only around 100 points (in our case) related to the human body, often missing data for specific body parts. Furthermore, the multi-path effect introduces ghost points, adding noise to the already sparse data. This sparsity and noise significantly challenge mmWave radar’s ability for scene flow estimation, as the lack of sufficient local geometric cues and the presence of noisy points make it difficult to extract robust local features, essential for accurately tracking movements within the scene.
Lack of Temporal Cues. The radar Doppler velocity measurement, representing the radial velocity of points, could enhance radar scene flow estimation. However, its resolution is limited by hardware and often too low to accurately capture small-scale movements typical in human sensing scenarios. Additionally, some radars, like the Vayyar radar [4] used here, do not measure Doppler velocity due to their specific signal transmitting design, further complicating accurate motion estimation. This limitation, coupled with the absence of consistent radar point data for certain body parts across frames, presents non-trivial challenges for reliable scene flow estimation as they result in the lack of temporal cues.
Scene Flow Annotation. Learning accurate scene flow estimation with deep networks requires point-level scene flow labels for training. However, manual annotation is prohibitively expensive and lacks real-world correspondence. Recent approaches [39, 35, 22] annotate object bounding boxes to generate pseudo scene flow labels, effective in autonomous driving with rigid bodies like cars. Nevertheless, this method is still labour-intensive and less applicable to human sensing due to the non-rigid nature of human movement, posing a challenge in creating accurate training labels for mmWave radar-based scene flow estimation.